OpenStudy (anonymous):

a fair coin is tossed 12 times. what is the probability of obtaining exactly 10 heads

OpenStudy (kropot72):

The probability of a head on each toss is 0.5. The probability of exactly 10 heads in 12 tosses is: \[P(10H\ out\ of\ 12\ tosses)=\left(\begin{matrix}12 \\ 10\end{matrix}\right)0.5^{10}0.5^{2}=66\times (0.5)^{12}\]